# Entropy Collapse in Mobile Sensors: The Hidden Risks of Sensor-Based Security

### v7

@@ Abstract
Mobile sensor data has been proposed for security-critical applications such as device pairing, proximity detection, and continuous authentication. However, the foundational premise that these signals provide sufficient entropy remains under-explored. In this work, we systematically analyse the entropy of mobile sensor data across four diverse datasets spanning multiple contexts. Our findings reveal pervasive biases, with single-sensor mean min-entropy values ranging from 3.408–4.483 bits ($\sigma$=1.018–1.574) despite Shannon entropy being several multiples higher, showing a significant collapse between average- to worst-case settings. We further demonstrate that correlations between sensor modalities reduce the worst-case entropy of using multiple sensors by up to $\approx$75compared to average-case Shannon entropy. This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits of min-entropy when combining 20 sensor modalities. Our results demonstrate that adversaries may feasibly predict sensor signals through an exhaustive, brute-force exploration of the entire measurement space. Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security, and we strongly urge caution when relying on mobile sensor data for security applications.
@@ 1 Introduction
Modern mobile devices are equipped with various embedded sensors—accelerometers, gyroscopes, magnetometers, and others—that capture continuous motion and environmental data at fine temporal granularity. This sensor data has enabled myriad applications from activity recognition to context-aware computing. More recently, research has proposed leveraging these signals for security-critical tasks such as cryptographic key generation, zero-interaction device pairing, and continuous authentication. A crucial yet under-explored assumption underpins such designs: sensor data provides sufficient unpredictability to thwart adversarial inference. Traditional “shake-to-pair” protocols [1] rely on motion patterns to establish secure communication between co-located devices, while other methods have incorporated ambient phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [2, 3, 4, 5, 6] and reduce user authentication prompts [7, 8, 9, 10, 11].
Despite these advancements, fundamental issues remain: multi-modal sensing is often advocated to counter sensor-specific weaknesses [2, 4, 12, 13], but the quantitative security benefits of combining multiple sensors has not been rigorously evaluated. Many existing studies rely on heuristic assessments or machine learning classifiers, e.g. [4, 13, 14, 2, 12], that do not address critical security questions. That is, firstly, how much entropy do sensors truly provide? And, secondly, to what extent do multi-modal sensor combinations provide security gains? Understanding the underlying entropy is important: even if different sensors are combined, fused or otherwise transformed, it does not fundamentally improve the quantity of entropy, or unpredictability, inherent in such signals. This paper investigates those concerns.
Our analysis reveals systemic limitations: commodity sensors exhibit significant biases of between 3.408–4.483 bits of min-entropy (5.584–9.266 bits of Shannon entropy on average). In this paper, we analyse 25 different sensors compared to a far smaller number explored in related work, i.e. [15] (1 sensor), [16] (10), [7] (2), and [17] (3). Furthermore, to the best of our knowledge, we also present the first multi-modal entropy analysis at this scale. We find that, while multi-modal sensor usage confers some benefits, non-uniform distributions and inter-sensor correlations significantly reduce the worst-case min-entropy by $\approx$40–75compared to average-case Shannon entropy. These findings challenge the notion that increasing the number of sensors reliably strengthens security, and it underscores the inadequacy of using sensor data as a dependable entropy source. Our contributions are as follows:
@@
    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate, or brute force, the space of possible measurements. This gives rise to fundamental security risks to schemes that rely on signals from single and multiple mobile sensors.
  • 4.
    Ultimately, we do not recommend relying on commodity sensors as a source of unpredictability for security-critical applications, both on a single- and multiple-sensor basis.
The rest of this paper is organised in the following way: §2 discusses sensor-based security mechanisms and established entropy metrics. §3 explains our experiment design for entropy estimation, including the threat model and dataset selection. §4 presents empirical results of our analyses and §5 discusses the implications for system design. We conclude in §6 with recommendations for further work. Our analysis work is released publicly to foster future research. [^2] https://github.com/cgshep/entropy-collapse-mobile-sensors
@@ 2 Background
This section discusses sensor-based applications, critiques existing approaches, and formalises the entropy metrics underpinning our analysis.
2.1 Mobile Sensor-based Security Applications
One major security application of mobile sensors is proximity detection, particularly for mitigating man-in-the-middle and relay attacks on mobile devices [18, 13, 14, 4, 2]. Mehrnezhad et al.[13] introduced a technique that uses accelerometer readings to verify that an NFC payment instrument and terminal are physically tapped together. The authors posit that “physical tapping causes random but correlated vibrations at both devices, which are hard to forge (or reproduce)” (p.1, [13]). The work reports an equal error rate (EER) of 17.65% using a machine learning-based approach. Gurulian et al. [14] also explored the use of shared vibration patterns between users using unique vibration patterns generated by one device. Shrestha et al. [2] explore four environmental modalities, including temperature and humidity, for proximity detection. In this work, single sensors yields 0.733–0.881 F1-score, while combining multiple increases the overall performance to 0.913–0.957. Mobile sensors have also found utility in tackling the longstanding problem of continuous authentication [19, 20, 21, 8, 9, 22, 10, 11], where sensor data is used to authenticate users passively without explicit interaction (with up to 99% accuracy claimed in some work [20]). Sensors have also been used to underpin the security of novel device pairing schemes, whether as a primary [23] or second line of authentication [1]. These schemes generally rely on detecting similar motion patterns between two devices using joint accelerometer and gyroscope sensor measurements in order to provide evidence of co-location.
The notion of ‘hardness’ is typically inferred through model evaluation metrics. A general approach follows one whereby sensor data is collected from $N$ users from which various features are extracted in the time or frequency domain (e.g. cross-correlation, spectral energy, Hamming, Euclidean and mean-absolute distances) [4, 13, 3, 2]. Features are then classified using simple threshold-based or supervised classification models, e.g. Support Vector Machines (SVM) and Random Forests. Calculating false positive (FPR) and negative (FNR) rates [1], precision and recall [8], EERs [14, 13], Receiver-Operator Curves (ROC) [11], and accuracy [11] are used to evaluate the model with respect to distinguishing between legitimate and illegitimate samples. A system is deemed to be effective if the model can discriminate between such samples with low error. (User studies have also been employed to evaluate the effectiveness of sensor-based authentication systems [21, 1].)
Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a minority of work in the literature. T2Pair by Li et al. [24], a zero-interaction pairing protocol, is found to have 32.3–38.5 bits of Shannon entropy; a refinement by Wu et al. [25] reported 51–54 bits. Even in the best cases, this is low relative to modern cryptographic standards. [^3] See AIS 20/31 [26]: FCS_RNG.1 specifies $\geq$240 and $\geq$250 bits of min- and Shannon entropy respectively for the effective internal state of a random number generator. We also point to work that has questioned the utility of mobile sensors in time-critical domains. Markantonakis et al. [4], building on Gurulian et al. [5] and Shepherd et al. [6], presented a reproducibility study of mobile sensors when deployed under a 500ms time constraint for NFC-based transactions as specified by the EMV payment protocol. Here, 0.179–0.246 EER was reported depending on the given sensor combination. Sensors were thus deemed unsuitable for proximity and relay attack detection without posing usability and security issues in practice.
2.2 Sensor Entropy Analyses
@@
In earlier work, Voris et al. [15] investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone. The authors find that min-entropy—defined in §2.3—is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy. Intrinsic noise from the sensor’s circuitry and seismic noise, and the sampling rate of its analog-to-digital converter (ADC), are considered significant influences on entropy generation. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. [17] analysed three mobile sensors on an undisclosed Xiaomi Redmi smartphone: a triaxial accelerometer, gyroscope, and magnetometer. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method. Krhovják et al. [7] examined the entropy of image and audio data collected from mobile phone cameras and microphones respectively. Using Nokia N73 and E-Ten X500 and M700 phones, Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera) are reported, with min-entropy of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. [16] presented an analysis of 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. A single-sensor analysis is presented, yielding min-entropy values of 0–7.85; motion sensors, e.g. accelerometer and vibration sensors, produced the highest entropy.
@@
Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices. Suciu et al. [27] proposed using a phone’s GPS module along with its accelerometer, gyroscope and orientation sensors. Using data from an HTC Google Nexus One, the approach passes the tests established in the NIST SP 800-22 [28] suite, but no precise entropy values were presented. Wallace et al. [29] explored an RNG design using the accelerometer, gyroscope, microphone, WiFi, GPS and camera data as randomness sources. Results are presented from a non-standard entropy evaluation using 37 Android devices. Sensors within existing work serve as inherently opportunistic entropy sources, i.e. in contrast with dedicated TRNGs using ring oscillators, Johnson-Nyquist thermal noise, and quantum phenomena (e.g. see [30]). Mobile sensors depend heavily on user behaviour and environmental, which could result in biases and correlations that are absent in controlled entropy sources. Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics [8, 11]. Such approaches do not account well for skewed distributions and other biases in the underlying data that affect predictability. Contrast this with typical measures used in the area of randomness testing and authentication [31, 32, 30, 33, 26]. For instance, NIST SP800-90B [33] and AIS 20/31 [26] recommend the use of min-entropy to estimate ‘worst-case’ unpredictability of a given source. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
@@ 2.3 Definitions
We use the following definitions and notation throughout this work.
@@
We draw attention to four special cases of $\alpha$ that are widely used in the literature. Firstly, the Hartley (Max) Entropy, given in Eq. 2, is the logarithm of the number of possible outcomes that have non-zero probability; it serves effectively as an upper bound.
@@
Second is the Shannon Entropy (Eq. 3), which corresponds to the classical definition of entropy in information theory, and is the limit of $H_{\alpha}$ as $\alpha\to 1$.
@@
Another case is the Collision Entropy ($\alpha=2$), quantifying the probability of “collisions” of multiple draws from $X$. This is given in Eq. 4.
@@
$H_{1}$ provides an average-case measure of uncertainty. It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage. Min-entropy is thus used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes. This is the Min-Entropy (Eq. 5), i.e. the value in the limit $\alpha\to\infty$.
@@
Note that $H_{\alpha}$ is a non-increasing function of $\alpha$, i.e. $H_{\infty}(X)\leq H_{2}(X)\leq H_{1}(X)\leq H_{0}(X)$. $H_{\infty}(X)$ focuses on the single most likely outcome, providing a strictly tighter (and generally minimum) bound on uncertainty. We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in NIST SP800-90B [33] and AIS 20/31 [26]. We also rely on joint entropy for assessing the entropy of multiple random variables.
@@
$H_{\alpha}(X_{1},X_{2},\dots,X_{n})\,=\,\frac{1}{1-\alpha}\,\log\Biggl{(}\sum_{(x_{1},\dots,x_{n})}p(x_{1},\dots,x_{n})^{\alpha}\Biggr{)}$ (6)
In the limit $\alpha\to 1$, then $H_{\alpha}(X_{1},X_{2},\dots,X_{n})$ converges to the classical joint Shannon entropy of these $n$ variables. To support the later discussion on Chow-Liu trees—our approach to joint entropy estimation—we also define the Kullback-Leibler divergence.

### v8

@@ Abstract
Mobile sensor data has been proposed for security-critical applications such as device pairing, proximity detection, and continuous authentication. However, the foundational premise that these signals provide sufficient entropy remains under-explored. In this work, we systematically analyse the entropy of mobile sensor data using four datasets from multiple application contexts (UCI-HAR, SHL, Relay, and PerilZIS). Using direct computation and estimation, we report entropy values—max, Shannon, collision, and min-entropy—for an exhaustive range of sensor combinations. We demonstrate that the entropy of mobile sensors remains far below what is considered secure by modern standards for security applications, such as for authentication and key generation systems, even when many sensors are combined. In particular, we observe an alarming divergence between average-case Shannon entropy and worst-case min-entropy. Single-sensor mean min-entropy varies between 3.408–4.483 bits despite Shannon entropy being several multiples higher. We further show that correlations between sensor modalities contribute to a $\approx$75% reduction between Shannon and min-entropy. This brings joint min-entropy well below 10 bits in many cases and, in the best case, yielding only $\approx$24 bits when combining 20+ sensor modalities. Our results reveal that adversaries may feasibly predict sensor signals through an exhaustive exploration of the measurement space. Our work also calls into question the widely held assumption that adding more sensors inherently yields higher security. Ultimately, we strongly urge caution when relying on mobile sensor data for security applications.
@@ 1 Introduction
Modern mobile devices expose rich motion and ambient signals from accelerometers, gyroscopes, magnetometers, and more. Beyond activity recognition and context inference, e.g. [1, 2, 3, 4], a long line of work proposes using these signals for security-critical tasks. This includes cryptographic key generation, zero-interaction device pairing, and continuous authentication [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Traditional key generation protocols [8, 18] have historically relied on shared motion patterns in order to establish secure communication between co-located devices. Meanwhile, other proposals have incorporated environmental phenomena, such as characteristics of magnetic fields and thermal fluctuations, to mitigate relay attacks [9, 10, 21, 5, 11] and to reduce the frequency of user authentication prompts [17, 16, 12, 13, 14].
Despite these advances, a crucial assumption remains under-examined: that sensor data provides sufficient entropy to thwart adversarial inference. Many existing studies rely on heuristic assessments or model classsification metrics, e.g. [21, 7, 6, 9, 22, 19, 20], that do not address two basic questions: (1) how much entropy do sensors actually provide, and (2) to what extent do multimodal sensor combinations improve entropy? For the latter, multimodal sensing has been advocated to counter sensor-specific weaknesses [9, 21, 22, 7], but the quantitative security benefits of combining multiple sensors has not been rigorously evaluated.
The security implications of low-entropy sensor data can be direct and severe. If such data is used as authentication evidence or to derive secret keys, an adversary may attempt to enumerate the entire space of possible measurement values in order to (re-)derive it. For such reasons, modern security standards provide important quantitative thresholds for what can be considered ‘secure’. For instance, NIST SP 800-57 [23] states that security levels below 80 bits are broken for long-term use, with 112 bits as the current minimum and 128 bits recommended for security keys. Standards that govern dedicated entropy sources are even stricter: NIST SP 800-90B [24] expects entropy sources used for cryptographic random number generators (RNGs) to approach full per-bit entropy, [^2] More specifically, $\approx$1 bit ($\epsilon\leq 2^{-32}$ per-bit) of min-entropy per output bit. while AIS 20/31 [25] requires an RNG’s internal state to possess over 240 and 250 bits of min-entropy and Shannon entropy respectively. As we will demonstrate, the min-entropy of even complex combinations of mobile sensors falls short of these standards.
In this paper, we analyse the entropy of 25 mobile sensors, which is substantially more than prior work—[26] (1 sensor), [27] (10), [17] (2), [28] (3)—across multiple public datasets. Our results reveal structural limits. Commodity sensors exhibit pronounced biases and dependencies, yielding on average 3.408–4.483 bits of min-entropy and 5.584–9.266 bits of Shannon entropy per sample across datasets. Multimodal fusion confers only modest gains: inter-sensor correlations and statistical bias both contribute to a divergence between average- and worst-case entropy, with min-entropy typically 40–75% below Shannon entropy for the same signals. In practical terms, schemes that treat these measurements as secrets expose a guess space small enough to be systematically enumerable. Our contributions are as follows:
@@
    We show how the collapse in worst-case entropy opens the door to attacks that exhaustively enumerate the space of possible measurements. This raises fundamental concerns with respect to security applications that rely on signals from single and multiple mobile sensors. Consequently, we do not recommend relying wholly on mobile sensors as entropy sources for security-critical applications.
The rest of this paper is organised as follows: §2 discusses sensor-based security mechanisms and established entropy metrics. §3 explains our experiment design for entropy estimation, including the threat model and dataset selection. §4 presents empirical results of our analyses and §5 discusses the implications for system design. We conclude in §6 with recommendations for further work. Our analysis work is released publicly to foster future research. [^3] https://github.com/cgshep/entropy-collapse-mobile-sensors
@@ 2 Background
This section discusses sensor-based applications and existing approaches, and formalises the entropy metrics underpinning our subsequent analyses.
2.1 Related Work
Mobile sensing has been widely explored for security tasks in three broad families: (1) proximity detection and relay-attack mitigation, where co-presence is inferred from shared motion or ambient context [29, 7, 11, 5, 21, 6, 9]; (2) device pairing, where two devices derive evidence of co-location to establish a secure association or shared secret [8, 30, 18, 20]; and (3) continuous authentication, which aims to verify users passively during normal interaction [16, 12, 13, 14, 31, 32, 33, 34]. Representative designs include tap-to-pair schemes using accelerometers [7], induced shared vibrations [6], and ambient-signal approaches that leverage temperature, humidity, magnetic field, or other environmental cues [9, 21]. These mechanisms commonly fuse accelerometer and gyroscope measurements to capture correlated motion as evidence of co-presence [8, 30].
Our key observation is that model-performance metrics are typically decoupled from the entropy of the underlying data sources upon which such models are based. Across the aforementioned application areas, effectiveness is usually inferred from discriminative performance. A common pipeline—documented across proximity detection, pairing, and continuous authentication—collects data from $N$ participants; derives time- or frequency-domain features (e.g. cross-correlation, spectral energy, and Hamming, Euclidean, and mean absolute distances); learns a decision rule (thresholds or supervised models such as SVMs, Random Forests, Naïve Bayes); and reports FPR/FNR, EER, ROC/AUC, accuracy, or F1. This template appears in numerous systems, including [7, 9, 21, 11, 16, 14, 13, 31, 32, 6]. For a comprehensive treatment of the model evaluation metrics used by existing mobile sensor-based security systems, the reader is referred to the surveys by Abuhamad et al. [19], Alzubaidi and Kalita [35], and Mayrhofer and Sigg [20]. While discriminative performance metrics are informative for usability and functionality, they do not quantify entropy of the underlying sources, which is the focus of our analysis.
Adjacent evidence also highlights practical limits under tight timing constraints. For NFC transactions, EMV specifies $\approx$500 ms for proximity checks. In that setting, Markantonakis et al. [21] (building on [5, 11]) report EERs of 0.179–0.246 across sensor combinations, concluding that commodity sensors are unsuitable for reliable proximity and relay detection without incurring usability or security issues. Taken together, the literature demonstrates functional feasibility but leaves open the core question of how unpredictable the underlying measurements really are, especially under multimodal fusion. This motivates our entropy-centric evaluation in the remainder of the paper.
Some proposals have attempted to assess the entropy of sensor signals within the context of a security mechanism, but this represents a small minority of the literature. Specifically, T2Pair by Li et al. [15] reports 32.3–38.5 bits of Shannon entropy, and a refinement by Wu et al. [36] reports 51–54 bits. Even in the best cases, these values are low relative to modern cryptographic standards; moreover, we are aware of no work using standard min-entropy as a worst-case measure of unpredictability [24, 25, 37] (see comparison tables of state-of-the-art schemes in [35, 18, 20]). We address this gap by analysing the entropy of a wide range of mobile sensors, both in isolation and in combination.
@@ 2.1 Related Work
Concurrent work has attempted to assess the entropy bestowed by mobile sensors as a separate line of inquiry. Voris et al. [26] investigated accelerometers as true random number generators (TRNGs) on a WISP RFID tag and Nokia N97 phone. The authors find that min-entropy is proportional to the motion applied to the device, with stationary movement having the lowest min-entropy. Intrinsic noise from the sensor’s circuitry and seismic activity, and the sampling rate of its analog-to-digital converter (ADC), are found to influence entropy generation. Min-entropy values of 3.1–11.4 bits were measured depending on the movement of the accelerometer. Lv et al. [28] analysed three mobile sensors—a triaxial accelerometer, gyroscope, and magnetometer—on an undisclosed Xiaomi Redmi smartphone. Min-entropy values of 0.593–5.876 are reported, depending on the modality and the entropy estimation method. Krhovják et al. [17] examined image and audio data collected from mobile phone cameras and microphones respectively on a Nokia N73 and E-Ten X500 and M700 phones. The work reports Shannon entropies of 2.9 (microphone) and 2.408–5.376 (camera), and min-entropies of 0.5 (microphone) and 0.754–3.928 (camera). Hennebert et al. [27] analysed 10 sensors on two wireless sensor monitors: a TI eZ430-RF2500 and a Zolertia Z1. A single-sensor analysis is presented, yielding min-entropy values of 0–7.85; motion sensors, e.g. accelerometer and vibration sensors, produced the highest entropy.
@@
Sensors have also been suggested as entropy sources in low-cost RNG designs for mobile devices. Suciu et al. [38] proposed combining GPS along with accelerometer, gyroscope and orientation sensors on an HTC Google Nexus One, passing NIST SP 800-22 tests [39], but without precise entropy estimates. Wallace et al. [40] explored a multi-sensor RNG design (accelerometer, gyroscope, microphone, WiFi, GPS, camera) with a non-standard entropy evaluation using 37 Android devices. Unlike dedicated TRNGs (e.g. ring oscillators, Johnson–Nyquist thermal noise, quantum sources [41]), mobile sensors depend heavily on user behaviour and environmental factors, introducing biases and correlations absent in controlled entropy sources. Existing sensor-based mechanisms largely overlook these dynamics, relying principally on heuristic or model evaluation metrics that do not account well for skewed distributions and other biases affecting predictability. In contrast, proposals of new entropy sources and authentication mechanisms commonly rely on worst-case metrics, such as min-entropy as stipulated in NIST SP 800-90B and AIS 20/31 [42, 43, 41, 24, 25]. Our study bridges the gap by systematically evaluating entropy across modalities and datasets.
Table 1: Comparative summary of prior sensor entropy analyses.
Reference Sensorsa Metricsb Results (bits) Context / Use-Case
Voris et al. [26] Acc. $H_{\infty}$ 3.1–11.4 True random number generation
Hennebert et al. [27] 10 (Various) $H_{\infty}$ 0–7.85 Harvesting entropy from physical mobile sensors
Lv et al. [28] Acc., Gyr., Mag. $H_{\infty}$ 0.593–5.876 Entropy source analysis on smartphone
Krhovják et al. [17] Cam., Mic. $H_{\{1,\infty\}}$ 0.5–5.376 Randomness sources in mobile devices
Li et al. [15] Mic., Light $H_{1}$ 32.3–38.5 Zero-interaction device pairing
Wu et al. [36] Mic., Light $H_{1}$ 51–54 Zero-interaction device pairing
This work 25 (Various) $H_{\{0,1,2,\infty\}}$ 8.1–23.9 ($H_{\infty}$) Systematic entropy analysis, including sensor combinatons, for security systems
  • a
    Acc.: Accelerometer; Gyr.: Gyroscope; Mag.: Magnetometer; Cam.: Camera; Mic.: Microphone.
  • b
    $H_{0}$: Max; $H_{1}$: Shannon; $H_{2}$: Collision; $H_{\infty}$: Min-entropy.
In general, while several studies have investigated the entropy of specific sensors, a direct comparison is often difficult due to varying methodologies and contexts. Table 1 provides a structured summary of key prior works, contrasting the sensors analysed, entropy metrics used, and the reported entropy ranges.
@@ 2.2 Definitions
We use the following definitions and notation throughout this work. Furthermore, unless stated otherwise, all logarithms are base 2 and entropies are reported in bits.
@@
We draw attention to four special cases of $\alpha$ that are widely used in the literature. First, the Hartley (Max) Entropy (Eq. 2) is the logarithm of the number of possible outcomes with non-zero probability and serves as an upper bound.
@@
Second, Shannon entropy (Eq. 3) is the classical definition of entropy in information theory, corresponding to the limit of $H_{\alpha}$ as $\alpha\to 1$.
@@
Third, the collision entropy ($\alpha=2$) quantifies the probability of drawing the same value twice from $X$:
@@
We note that $H_{1}$ provides an average-case measure of uncertainty. It takes into account the entire distribution of outcomes; however, an adversary may only need to guess the most likely event to gain an advantage. Min-entropy (Eq. 5) is used as a conservative, worse-case metric, accounting for the least favorable distribution of outcomes; it is the value in the limit $\alpha\to\infty$:
@@
$H_{\alpha}$ is a non-increasing function of $\alpha$, i.e. $H_{\infty}(X)\leq\dots\leq H_{1}(X)\leq H_{0}(X)$. $H_{\infty}(X)$ focuses on the single most likely outcome, providing a strictly tighter (and generally minimum) bound on uncertainty. We observe that min-entropy ensures that even the most skewed probability distributions still meet the required security guarantees; indeed, it is a recommended method for assessing entropy sources in modern standards [24, 25]. We also rely on joint entropy for assessing the entropy of multiple random variables.
@@
$H_{\alpha}(X_{1},X_{2},\dots,X_{n})\,=\,\frac{1}{1-\alpha}\,\log\Biggl{(}\sum_{(x_{1},x_{2},\dots,x_{n})}p(x_{1},x_{2},\dots,x_{n})^{\alpha}\Biggr{)}$ (6)
In the limit $\alpha\to 1$, then $H_{\alpha}(X_{1},X_{2},\dots,X_{n})$ converges to the classical joint Shannon entropy of these $n$ variables. To support the later discussion on Chow-Liu trees—our approach to joint entropy estimation—we define the Kullback-Leibler divergence.